solve the system of equations by graphing.
3x-18=9y
2x-6y=-12

Answer by mananth(16946)

3x-18=9y
/9
1/3 x -2=y.............................(1)
2x-6y=-12
-6y=2x-12
/-6
y=1/3 x+2............................(2)


equation1
when x= 0 y= -2 ( 0 , -2 )
when y = 0 x= 6 ( 6 , 0 )
Equation 2
when x= 0 y= 2 ( 0 , 2 )
when y = 0 x= -6 ( -6 , 0 )

The lines are parallel hence NO SOLUTION
m.ananth@hotmail.ca